# Tree networks of real-world data: analysis of efficiency and spatiotemporal scales

### v3

@@ 5 Results
In panels (a)-(d) the pattern-matching efficiency, $\theta$, is depicted. The vertical red-dashed line indicates the $r$ value for the optimal trade-off, while the shaded purple region indicates a $r$-domain tolerance. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
@@
In panels (e)-(h) the thermodynamic efficiency, $\eta$, is depicted and, for the case of the NIST and the FashionMNIST a notable consistency in the global maxima is observed with the pattern-matching efficiency $\theta$ in terms of optimal $r$. See the $r$-range spanned by the purple-shaded areas in panels (a)-(d) and the vertical dashed lines in panels (e)-(h). This consistency in the maxima of the two efficiencies suggests that thermodynamic or pattern-matching might be related to the context of this algorithm. In the Section 6 we discuss this matter.
@@
Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph. Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset. Such transition is also depicted in Figure 7 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 6. See also Section S4 illustrating the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
@@
Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Section 4. Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (i)-(l) of Figure 6 and in panels (a)-(c) of Figure 7. In the seminal work [1], it is indicated to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
@@
Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. As shown in panel (h) of Figure 5 and in panels (a)-(d) of Figure 7 where the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves are illustrated. The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat. Thus, it is crucial to identify which is the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in panel (b) of Figure 5. Such resolution scale is the scale we consider of significant importance for the computation of the efficiency $\eta$ across the transition and defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In Section S3 the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
@@
Moreover note in Figure 7 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition. We note that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
@@
Figure 6: Thermodynamic & pattern-matching efficiency. In (a)-(d) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In (e)-(h) the optimal thermodynamic efficiency, $\eta$, as a function of $r$. In panels (f) and (h) a noticeable cohesion of the global maxima of the two efficiencies is observed. Vertical dashed lines indicate the global maxima of the efficiencies. Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure. Note that for datasets of higher complexity the highest efficiency, both thermodynamic and pattern-matching, tends to be posited closer to the structural phase transition to scale-invariant illustrated by the specific heat in panels (i)-(l). This is in line with the argument that states that a fairly simple dataset does not benefit from the vicinity of the phase transition. Quantities are averaged over 50 independent simulations.
@@ 5.1 Performance at the emergence of scale-invariance
As previously discussed in Section 4, in panels (i)-(l) of Figure 6 and in panels (a)-(d) of Figure 7, the $\tau$-scale where the bifurcation shows up indicates the scale involved in the structural transition, and it is the resolution scale where the thermodynamic efficiency for some of the datasets displays consistency with the pattern-matching efficiency as a function of $r$. Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
To discuss this empirical result we turn our attention to the panels (a)-(d) and (i)-(l) in Figure 6 and panels (a)-(d) in Figure 7. For the CIFAR10 and FashionMNIST datasets there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the global maximum of the efficiencies is observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively for the two datasets. On the other hand, for the MNIST and NIST datasets the thermodynamic and pattern-matching efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$. The term "far away" is quantified by the purple-shaded areas in panels (a)-(d) of Figure 6. The argument also holds for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area. For the MNIST case, we observe that a third large peak emerges in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 6 and the entropy curves depicted in Figure S6). This effect is only mitigated for larger datasets, i.e. larger networks, as we show in Section S2 of the Supplementary. This intermediate peak is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$. Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 6(c), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 6(g). We stress that also in this case, we measured $\eta$ at the $\tau$ scale indicated by the bifurcation $\tau\sim 10$.
@@
To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or farther away, we turn to the argument, already present in the literature [37, 10, 38], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only if a high level of complexity is required by the task, see Figure 2. While for the CIFAR10 and FashionMNIST, more complex datasets when compared to both the MNIST and NIST, their respective optimal value for the efficiency is posited in the proximity of the phase transition, the optimal values of $r$ for the NIST and MNIST datasets are far away from their respective transition values. The MNIST, that is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$. While the global maximum in Figure 6(g) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 7, the local maximum is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation. This observation is in line with the hypothesis that the system takes advantage of the critical behavior of a certain resolution region only when a fairly complex task is presented to it and in line with other observations discussed in the literature [37, 10]. Thus, it stresses the importance of the dataset under analysis and its disordered distribution of points in a high-dimensional feature space.
[…]
@@ 6 Discussion
The ability of a system to survive is the capability of concentrating a flow of order from the environment - “drinking order from the environment” [39]. In other words, systems, whether living systems or machine(-learning) systems, need to learn an efficient internal representation of the complex world - or the dataset- they are posed in contact with. Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or in the very network-features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [40], or in a mixed-way as in artificial neural networks where the engineer designs a complicated distribution through the definition of an architecture and the training phase fills in the parameters in order to obtain the desired function-approximation machine to achieve statistical generalization [41].
@@
When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42]. Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g. class-matching.
@@
The recently developed theoretical tools of statistical physics of fields for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
@@
In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure. The tree-formation undergoes two phase transitions at two different scales: a meso-scale with the emergence of a approximately scale-invariant structural behaviour that progressively involves a larger scales, and a large-scale transition from the irregular scale-invariant behaviour to the regular-scale invariant 1d lattice. Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different level of complexity. When the dataset was more difficult, an optimal efficiency to the pattern-searching task is found to be close to the transition to scale-invariant in the parameter space. Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range. Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an undirected process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
@@
In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$. For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$. It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4]. They are not exactly alike as one indeed involves a directed process while the former involves an undirected process; still, they can be used as estimates of the "flow".
@@
We have discussed the denominator of the two efficiencies: $\eta=U/W$ and $\theta=\text{Accuracy}/\log N_{dist}$. We now discuss the numerator. For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se as evidenced by the fact that the accuracy is not necessarily a monotonic decreasing function of the control parameter $r$, but it can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.
@@
The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping. Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one to one relation between the two. In the case of a network structure the parameters are the network traits themselves, which in the Fourier space become the “network frequencies” $\lambda_{i}$, i.e. the Laplacian eigenvalues. We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5. The internal energy $U$ is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale $\tau$, tantamount to the inverse temperature $\beta$ in statistical physics. Thus, the Gibbs average frequency content is quantified by $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$ the internal energy, where for scale-invariant networks $<\lambda>_{\tau}=c_{0}/\tau$ [3]. The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given an external control parameter $r$.
@@
Notably, we have computed the thermodynamic efficiency at a specific resolution scale $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an "ultra-violet" cut-off delimiting the relevant frequencies of the network structure across the transition to scale-invariance, i.e. from star graph to amorphous. The operational $\tau$-scale employed roughly identifies modules in the network as at the $r$-bifurcation point we have observed a phase transition when employing the modularity as an order parameter. While the transition point to scale-invariant permits the identification of the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency. We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
@@
In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of solely the Portegys algorithm. Given a pre-trained tree network structure on a dataset that is unknown, it is possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the thermodynamic efficiency on the phase portrait without performing any test phase. As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, it posits itself close to the phase transition boundary at the cost of a slower pattern search. Interestingly, our analysis on hierarchical tree networks could be also applied, if further developed, to other topics in artificial intelligence. We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing. In the context of pure neuroscience, it would also be interesting to design experiments where the dendritic structure is related to the capability of the neuron to deal with complicated input-output mapping of signals. Moreover, potential applications include the topic of neuromorphic computing [45, 46, 47] currently looking for a possible solution to the ever-increasing computational power demand of recent years. In the neuromorphic paradigm computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio. The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [48]. In materio tree-like structures are currently physically implemented. We mention tree structures produced at the critical point of a percolation transition [49, 50, 51] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 52], thus, their capabilities could be investigated and enhanced based on this work and future developments.

### v4

@@ 5 Results
Panels (a)-(d) depict the thermodynamic efficiency, $\eta$, and panels (e)-(h) depict the pattern-matching efficiency, $\theta$. The vertical dashed lines indicate the peaks of the efficiencies, while the shaded purple regions in (e)-(h) indicate a $r$-domain tolerance for $\theta$. The tolerance domain is defined as those $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
@@
Note the consistency in the global maxima of $\theta$ and $\eta$ in terms of optimal $r$ for the NIST and the FashionMNIST. See the $r$-range spanned by the purple-shaded areas in panels (e)-(h) and the vertical dashed lines in panels (a)-(d). This consistency in the maxima suggests that the two efficiencies might be related, for the specific context of this algorithm. In the Section 6 we discuss this matter.
@@
The topological complexity of the networks increases at a critical $r$ from the trivial traits of a star graph to the fractal structure of the amorphous network. Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=C_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset. Such transition is also depicted in Figure 8 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 7. See also Section S4 which illustrates the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
We note that the specific heat of the MNIST has a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 7). This is an intrinsic characteristic of the dataset, as it is only mitigated for larger networks, as shown in Section S2 of the Supplementary. This intermediate peak is sufficiently strong to be reflected on a local maximum of $\eta$, in Figure 7.(c). Although this local maximum in $\eta$ is not detected by $\theta$ in Figure 7(g), it is reflected in its standard deviation, allowing us to include the $r$-range indicated by the shaded purple area when we compare $\eta$ and $\theta$ peaks in terms of $r$. This shaded region corresponds to the $\eta$-peak-to-peak $r$-range in Figure 7.(c).
@@
Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Sections 4 and 6. Such critical resolution scale is identified by the bifurcation of the specific heat peaks in panels (i)-(l) of Figure 7 and in panels (a)-(d) of Figure 8. In the seminal work [1], it is suggested to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, and logarithmically divide $\tau_{diff}$ to define the large propagation scales, the middle scales, and the small scales. In fact, this eigenvalue should provide the time scale where the diffusion process has reached equilibrium.
Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. Panels (a)-(d) of Figure 8 depict the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves. The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat.
@@
It is crucial to identify the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in Figure 5.(b). We consider such resolution scale of significant importance for the computation of the efficiency $\eta$ across the transition as it defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In Section S3 of the Supplementary Material the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
@@
Finally, note in Figure 8 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$. We remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only, and the latter shapes the degree distribution only, as the number of nodes and edges is fixed across the transition. We note that the Portegy’s algorithm stays in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
@@
Figure 7: Thermodynamic & pattern-matching efficiency. In (a)-(d) the thermodynamic efficiency, $\eta$, as a function of $r$. The thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant "length" scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=C_{0}>0$, reflecting a roughly scale-invariant structure. In (e)-(h) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In panels (b)-(f) and (d)-(h) a noticeable cohesion of the peaks of the two efficiencies is observed. Vertical dashed lines indicate the global maxima of the efficiencies. The efficiency peaks, both $\eta$ and $\theta$, are always in the amorphous region of the phase plane. For datasets of higher complexity, they tend to be closer to the structural phase transition border separating the star-graph to the amorphous region. This agrees with the argument that a fairly simple dataset does not benefit from the vicinity of the phase transition. Quantities are averaged over 50 independent simulations.
@@ 5.1 Performance at the emergence of scale-invariance
In this section, we discuss the relation between the location of the pattern matching efficiency, $\theta$, on the phase plane of the algorithm and the complexity of the dataset, see Figure 2. We turn our attention to the panels (e)-(h) and (i)-(l) in Figure 7 and panels (a)-(d) in Figure 8.
For the CIFAR10 and FashionMNIST datasets, there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the peaks of each of the two $\theta$ efficiencies are observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively. Note that for the FashionMNIST, also the thermodynamic $\eta$ peaks in the same range of $r$.
On the other hand, for the MNIST and NIST datasets, $\theta$ peaks in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$. The term "far away" is quantified by the purple-shaded areas in panels (e)-(h) of Figure 7. Note the correspondence of the peaks of $\eta$ and $\theta$ for the two datasets in terms of $r$. The latter observation on the correspondence of $\theta$ and $\eta$ also holds for the MNIST, if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
To explain the varying locations of peak efficiencies, whether near the transition border or farther from it, we refer to an argument in the literature [37, 10, 38]. A system needs tuning, or self-tuning, at the edge of a phase transition only when the task demands a high level of complexity, specifically, this was investigated in the context of second-order phase transitions [37, 10].
@@
For the more complex datasets CIFAR10 and FashionMNIST, the optimal $\theta$ values are closer to the transition border. In contrast, the optimal values for the simpler NIST and MNIST datasets are located farther from their respective transition points. This observation also holds for the thermodynamic $\eta$ for the FashionMNIST, MNIST, and NIST. Moreover, the MNIST, an intermediate dataset in terms of complexity, is an interesting case. It shows two maxima for the thermodynamic efficiency, $\eta$. The global maximum in Figure 7(c) is far away from the critical value of $r$ illustrated by the bifurcation in panel (c) of Figure 8, while the local maximum is at $r\sim 0.35-0.40$, which corresponds to the critical value of $r$ indicated by the bifurcation.
These observations align with the hypothesis that the system exploits its proximity to the phase transition boundary only when presented with a relatively complex task. The overall discussion highlights the importance of the dataset under analysis, particularly its disordered distribution of points within a high-dimensional feature space.
Finally, we observe that while the value of $r$ where a structural transition appears is dependent on the dataset, i.e. $r=0.50$ for the NIST, $r=0.40$ for the MNIST, $r=0.35$ for the FashionMNIST and $r=0.50$ for the CIFAR10, the $\tau\sim 10^{1}$ scale is roughly the same for all the four datasets. Interestingly, the bifurcation $\tau$-scale is also unaffected by the number of nodes in the network (see Supplementary Section S2). We speculate about the possibility that this is due to the constraint on sparsity as it is an invariant feature over the cases investigated in this work.
@@ 6 Discussion
The ability of a system to survive is the capability of concentrating a flow of order from the environment - “drinking order from the environment” [39]. In other words, systems, whether living systems or machine(-learning) systems, need to learn an efficient internal representation of the complex world - or the dataset - they are posed in contact with. Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or into the very network features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [40], or in a mixed way as in artificial neural networks where the engineer designs a complicated distribution through the definition of the architecture and the training phase fills in the parameters to obtain the desired function-approximation machine to achieve statistical generalization [41].
@@
When confronted with the real-world environment, to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42]. Moreover, the same system needs to cope with the cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine learning, the cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g. class-matching.
@@
The recently developed theoretical tools of statistical physics for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure. The tree formation undergoes two phase transitions at two different scales: the former is a mesoscale transition with the emergence of an approximately scale-invariant structural behavior progressively involving larger scales; the latter is a large-scale transition from the irregular scale-invariant structure to the regular-scale invariant 1d lattice.
Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different levels of complexity. The optimal pattern-matching efficiency was always found in the amorphous regime, Nevertheless, when the dataset was more difficult, an optimal efficiency to the pattern-searching task was found close to the transition border separating the star-graph and the scale-invariant regimes.
@@
Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range. Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an ergodic process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
@@
In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$. For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$. It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4]. They are not exactly alike as only the latter is an ergodic process; still, they can be used as estimates of the "flow".
@@
We have discussed the denominator of the two efficiencies: $\eta$ and $\theta$. We now discuss the numerator. For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se as evidenced by the fact that the accuracy is not necessarily a monotonic decreasing function of the control parameter $r$, but it can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.
@@
The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping. Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one-to-one relation between the two. In the case of a network structure, the parameters are the network traits themselves, which in the Fourier space become the “network frequencies”, $\lambda_{i}$, i.e. the Laplacian eigenvalues. We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5. The internal energy, $U$, is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale, $\tau$, tantamount to the inverse temperature, $\beta$, in statistical physics. Thus, the Gibbs average frequency content is quantified by the internal energy, $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$, where for scale-invariant networks $<\lambda>_{\tau}=C_{0}/\tau$ [3]. The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given the value of the external control parameter, $r$.
@@
Notably, we have computed the thermodynamic efficiency at a specific resolution scale, $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an "ultra-violet" cut-off delimiting the relevant frequencies of the network structure across the transition to scale-invariance, i.e. from star graph to amorphous. The operational $\tau$-scale employed roughly identifies the scale where modules emerge in the network. For this reason, we could gauge the phase transition employing the network modularity, $Q$, as an order parameter. While the transition to scale-invariant permits the identification of the network resolution scale, $\tau$, where the thermodynamic efficiency needs to be measured to observe a peak, the corresponding transition value of the control parameter, $r$, does not always coincide with the global maximum in the efficiency. We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
@@
In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of only the Portegys algorithm. Given a pre-trained tree network based upon an unknown dataset, it might be possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the peaks of the pattern-matching and thermodynamic efficiency on the phase plane without performing any test phase. As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, the peak of the pattern-matching efficiency might be at the phase transition boundary providing higher accuracy at the cost of a slower pattern search. Interestingly, our analysis of hierarchical tree networks could be applied, if further developed, to other topics in artificial intelligence. We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing. In the context of pure neuroscience, it would also be interesting to design experiments where the dendritic structure is related to the capability of the neuron to deal with complicated input-output mapping of signals. Moreover, potential applications include the topic of neuromorphic computing [45, 46, 47] currently looking for a possible solution to the ever-increasing computational power demand of recent years. In the neuromorphic paradigm, computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio. The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [48]. Tree-like structures are currently physically implemented in materio. We mention tree structures produced at the critical point of a percolation transition [49, 50] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 51], but also fractal trees produced in the dielectric breakdown [52], thus, their capabilities could be investigated and enhanced based on this work and future developments.